OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form.

OpenStudy (karatechopper):

all right so do you know how to find the slope?

OpenStudy (anonymous):

y = 3x – 6 y = 3x – 3 y = 3x + 6 y = 3x + 3 one of these

OpenStudy (wasiqss):

y-y1=m(x-x1) m=(y2-y1)/(x2-x1)

OpenStudy (karatechopper):

well whats your slope?

OpenStudy (anonymous):

-9/-3

OpenStudy (karatechopper):

ok now simplify that

OpenStudy (anonymous):

-3/-1

OpenStudy (karatechopper):

and simplify that

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

3? or -3?

OpenStudy (karatechopper):

-3/-1= 3

OpenStudy (karatechopper):

therefore 3 is your slope..

OpenStudy (karatechopper):

now do you know what point slope form is?

OpenStudy (anonymous):

no

OpenStudy (karatechopper):

do you know how to write slope intercept form?

OpenStudy (anonymous):

no

OpenStudy (karatechopper):

then how can you have a question and you have no clue what the main part of it means? is this your homework?

OpenStudy (karatechopper):

o well

OpenStudy (karatechopper):

slope intercept form is y=mx+b y= what you solved for m=slope b= y intercept

OpenStudy (karatechopper):

y can also be the y coordinate